### v2

# Vortices and Factorization

Abstract
We review applications of factorization methods to the problem of finding stationary point vortex patterns in two-dimensional fluid mechanics. Then we present a new class of patterns related to periodic analogs of Schrodinger operators from the “even” bi-spectral family. We also show that patterns related to rational and soliton solutions of the KdV hierarchy constitute complete solution of the problem for certain classes of vortex systems.
1 Introduction, Stationary Vortex Configurations
In fluid mechanics, factorization methods found their applications in the theory of two-dimensional flows in two opposite limits: the infinite-viscosity limit of the free-boundary flows (see e.g. [26], [27] and references therein) and in the zero-viscosity limit. Here we review the latter, namely applications to the problem of finding classes of two-dimensional stationary vortex configurations in inviscid 2d flows.
Systems of point vortices are weak solutions of the two-dimensional Euler equation. Work on their classification started more than a century and a half ago and has been developed substantially over the past several decades (for general overview see e.g. [3, 4, 14, 31, 32] and references therein).
Motion of two-dimensional inviscid and incompressible liquid is described by the system of PDEs
$\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\nabla)\mathbf{v}=-\frac{\nabla P}{\rho},\quad\rho={\rm const},$ (1)
$\nabla\mathbf{v}=0,$ (2)
where ${\mathbf{v}}=\mathbf{v}(x,y,t)$, $\mathbf{v}=(u,w)$ is the two-dimensional velocity of the flow, and $P=P(x,y,t)$ stands for the pressure. Equation (1) is the Euler’s equation, while (2) is the continuity equation. In two dimensions, the vorticity, which is a curl of velocity, has a single component $\omega$:
$\omega=\frac{\partial w}{\partial x}-\frac{\partial u}{\partial y}.$
Taking curl of (1) we get
$\frac{d\omega}{dt}=\frac{\partial\omega}{\partial t}+(\mathbf{v}\nabla)\omega=0.$
[…]
$\omega=\sum_{i=1}^{N}{\cal Q}_{i}\delta(x-x_{i}(t))\delta(y-y_{i}(t))$
the strength (i.e. circulation) of each vortex ${\cal Q}_{i}$ is time independent and speed of $i$th vortex $(dx_{i}/dt,dy_{i}/dt)$ equals velocity of the flow averaged over a small circular domain centered at $(x_{i},y_{i})$. Since vorticity is a curl of velocity, for the system of point vortices we have
$u-{\rm i}w=U-{\rm i}W+\frac{1}{2\pi{\rm i}}\sum_{i=1}^{N}\frac{{\cal Q}_{i}}{z-z_{i}(t)},$ (3)
where we used complex notations $z=x+{\rm i}y$. In (3) the full velocity field $\mathbf{v}=(u,w)$ is a sum of irrotational (i.e. zero-curl) background flow $\mathbf{V}=(U,W)$ and the flow created by superposition of $N$ point vortices. By (2, 3) $\mathbf{V}$ is a zero-divergence field. Since it is irrotational, it is also a gradient of a potential and as a consequence the potential is a harmonic function. In sections 1-8 we will consider unrestricted flows in the whole plane with velocities bounded at infinity. The only harmonic functions satisfying above conditions are linear functions and the background flow has constant velocity (uniform background flow). Using (3), we find the average velocity of the flow in a small disc centered at $z=z_{i}$, thus, obtaining the following equations of motion of vortices in the uniform background flow
$2\pi{\rm i}\frac{d\bar{z}_{i}}{dt}=k+\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}},\quad i=1\dots N,$ (4)
where $k$ is a complex constant.
Classification of the stationary patterns of vortices is then reduced to the study of solution of system of $N$ algebraic equations
$k+\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}=0,\quad i=1\dots N.$ (5)
This system also has the 2d electrostatic interpretation: (5) is nothing, but an equilibrium condition of $N$ point electric charges ${\cal Q}_{1},\dots,{\cal Q}_{N}$ interacting pair-wise through 2d Coulomb (logarithmic) potential and placed in the homogeneous electric field $\bar{k}$. The equilibrium configuration is a stationary point of the electrostatic energy
$E=k\sum_{i}{\cal Q}_{i}z_{i}+\sum_{i<j}{\cal Q}_{j}{\cal Q}_{j}\log(z_{i}-z_{j})+{\rm c.c.},$
where “c.c.” stands for the complex conjugate.
The stationary patterns with $k\not=0$ are called translating equilibria (or translating configurations/patterns) because in the reference frame where the fluid rests at infinity, the whole pattern translates with constant velocity $(\text{Im}\,k,\text{Re}\,k)/(2\pi)$. Below we mainly deal with equilibria without background flow, i.e. $k=0$ equilibrium configurations (“static configurations/patterns”):
$\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}=0,\quad i=1\dots N.$ (6)
This system of equations is invariant under rigid motions and scaling of the plane. From invariance of stationary value of electrostatic energy under scaling of the plane it follows that the system has solutions only if $\sum_{i<j}Q_{i}Q_{j}=0$. O’Neil showed [31] that for almost all values of $Q_{i}$, satisfying the above condition, there are exactly $(N-2)!$ distinct configurations, i.e. modulo rigid motions and scaling of the plane, distinct solutions of (6).
In the case $k\not=0$ necessary and sufficient condition of existence of solutions is $\sum_{i}{\cal Q}_{i}=0$. For almost all values of ${\cal Q}_{i}$ satisfying this condition number of distinct, modulo translations of the plane, solutions of (5) is $(N-1)!$.
There are, however, “resonant” values of ${\cal Q}_{i}$, where number of configurations is infinite. This paper mainly deals with such resonant cases. Sections 10 and 11 are devoted to periodic (“vortex street”) versions of such systems.
We also briefly consider non-resonant cases in relationship with factorization method. These are equilibrium patterns in non-uniform background flows from section 9.
2 Vortices, Polynomials and Rational Primitives
Let us consider static equilibria of $N=l+m$ charges (or vortices) of two species. Without lost of generality we can set values of first $l$ charges to -1 and those of remaining $m$ charges to $\Lambda$:
${\cal Q}_{i}=\left\{\begin{array}[]{ll}-1,&i=1\dots l\\ \Lambda,&i=l+1\dots l+m\end{array}\right..$
Introducing polynomials in $z$
$p(z)=\prod_{i=1}^{l}(z-z_{i}),\quad q(z)=\prod_{i=1}^{m}(z-z_{l+i}),$ (7)
we first consider a situation where all charges have distinct positions, so $p$ and $q$ do not have multiple or common roots. One can check that the static equilibrium condition (6) for this system of two species is equivalent to the following bi-linear differential equation for the above polynomials
$p^{\prime\prime}q-2\Lambda p^{\prime}q^{\prime}+\Lambda^{2}pq^{\prime\prime}=0,$ (8)
where prime denotes differentiation wrt $z$. The equivalence can be established by the partial fraction decomposition of the ratio of (8) and $pq$.
Note, that (8) holds not only when $p$ and $q$ do not have common or multiple roots, but also in special cases when such roots exist. It trivially follows from the partial fraction decomposition of the ratio of (8) and $pq$ that if a common root exists, then it cannot be the simple one in both $p$ and $q$: Multiplicities $M$ and $M^{\prime}$ of this root, in $p$ and $q$ respectively, are constrained by the following relation:
$(M-\Lambda M^{\prime})^{2}=M+\Lambda^{2}M^{\prime}.$
In situation when a common root emerges as a limit of a generic solution, when $M$ vortices of strength $-1$ and $M^{\prime}$ vortices of strength $\Lambda$ collapse into a single vortex of strength $\Lambda M^{\prime}-M$, the limiting vortex patterns are called the collapse configurations. There are also exist common root solutions that are not collapse solutions. For example, Demina and Kudryashov [16] presented a family of such solutions using the fact that if a pair $p$, $q$ solves (8) for some $\Lambda$, then another pair $pq^{\Lambda(\Lambda-1)/2}$, $q^{\Lambda(\Lambda+1)/2}$, that corresponds to the same vortex pattern as the first one, satisfies
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}=0.$ (9)
The above equation, which is the $\Lambda=1$ specification of (8), is called the Tkachenko equation [38] in literature on the vortex patterns [^1] Eq. (9) first appeared in work of Burchnall and Chaundy [7] in different context.. Its generic (i.e. no-common root) solutions describe equilibria of two species of vortices (charges) of equal magnitude but opposite sign. Bartman [9] identified the Tkachenko equation with the recurrence relation (16) for the Adler-Moser polynomials. As follows from works by Burchnall and Chaundy [7] (see the next Section) any generic polynomial solution of the Tkachenko equation is a pair of the Adler-Moser polynomials [1]
$p(z)=P_{n-1}(z),\quad q(z)=P_{n}(z).$ (10)
In this way we get complete classification of static configuration of system of vortices (charges) of strengths ±$\pm 1$. The degree of the $n$th Adler-Moser polynomial equals $n(n+1)/2$. Therefore, for this system equilibrium is only possible if the number of negative and positive charges equals $l=n(n-1)/2$ and $m=n(n+1)/2$ respectively. The equilibrium coordinates of charges $z_{i}$ of each specie are roots of consecutive Adler-Moser polynomials. The $n$th Adler-Moser polynomial $P_{n}$ is also a function of $n$ free parameters, and the space of all possible equilibria of $N=l+m=n^{2}$ charges is a $n$-dimensional complex subspace of the $n^{2}$ dimensional complex space.
Details on Adler-Moser polynomials will be given in the following sections. Here, we note that these polynomials were first found by Burchnall and Chaundy [7], later re-discovered in context of the Hadamard problem [23] by Lagnese and Stellmacher, and then by Adler and Moser as polynomial $\tau$-functions corresponding to rational solutions of the KdV hierarchy of integrable PDEs [1]. Burchnall and Chaundy [7] constructed these polynomials as solutions of the Tkachenko equation (9) by considering it as a linear second order ODE for $p$, where coefficients of ODE are determined by some polynomial $P_{n}=q$. Then the requirement that its two linearly independent solutions $p=P_{n-1}$ and $p=P_{n+1}$ are both polynomial is equivalent to rationality in $z$ of two indefinite integrals (rational primitives)
[…]
It turned out that, apart from the Adler-Moser case there is another case with $\Lambda=2$ when there exists an infinite chain of polynomial solutions (8) that can be constructed with help of rational primitives [9, 24].
3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
Let us return to the generic $\Lambda$ case, and consider the bi-linear equation (8) as a second order linear ODE for $p=p_{n-1}(z)$ with coefficients of the equation being defined by $q=q_{n}(z)$. By elementary methods we find that the second linearly independent solution $p=p_{n}(z)$ equals
[…]
In order to build a recursive chain we require that both $p_{n-1}$ and $p_{n}$ be polynomial in $z$, so the primitive
[…]
must be rational. Alternatively, considering (8) as an ODE for $q$, we come to the requirement of rationality of
[…]
It is clear that pair of the above rationality conditions requires that both $2\Lambda$ and $2/\Lambda$ be integer, which is possible only if
$\Lambda\in\left\{\frac{1}{2},1,2\right\}.$ (13)
The case $\Lambda=1$ has already been mentioned and corresponds to the Adler-Moser polynomials. Since equation (8) is invariant under the involution $\Lambda\leftrightarrow 1/\Lambda$, $p\leftrightarrow q$, the cases $\Lambda=1/2$ and $\Lambda=2$ are equivalent, and, without loss of generality, both can be represented by the $\Lambda=2$ case.
Now, let us take $p$ and $q$ that do not have common/multiple roots and factorize $p(z)$ as $p(z)=(z-z_{i})\tilde{p}(z)$, where $z_{i},i\leq l$ is a root of $p$ (see definition (7)). Rationality of primitive (11) implies an absence of a simple pole at $z=z_{i}$, which is equivalent to the condition
$\frac{\partial}{\partial z}\left(\frac{q(z)^{2\Lambda}}{\tilde{p}(z)^{2}}\right)_{z=z_{i}}=0.$
Since $\tilde{p}(z_{i})=p^{\prime}(z_{i})$ and $\tilde{p}^{\prime}(z_{i})=p^{\prime\prime}(z_{i})/2$, we get the following equation
$p^{\prime\prime}(z_{i})q(z_{i})-2\Lambda p^{\prime}(z_{i})q^{\prime}(z_{i})=0$
which is clearly satisfied by solutions of (8). Similarly, rationality of (12) also follows from (8). It is worth mentioning that the converse statement also holds, i.e. the bilinear equation (8) follows from rationality of (11) and (12) (for details see [24]).
Thus, once a pair of polynomials $p,q$ without multiple and common roots satisfying (8,13) is found, one can build an infinite chain of polynomial solutions of (8) with the help of recursive procedure for $p_{n}$ and $q_{n}$ mentioned at the beginning of this section.
In more detail: Since $p=z^{l}+\dots$ and $q=z^{m}+\dots$ from the leading term of (8) we get the diophantine equation connecting $l$ and $m$
  • $\Lambda=1$ case: $(l-m)^{2}=l+m$,
  • $\Lambda=2$ case: $(l-2m)^{2}=l+4m$.
Its solutions are
$l=l_{i},\quad m=m_{i},\qquad{\rm or}\qquad l=l_{i},\quad m=m_{i+1},\quad i\in\mathbb{Z},$ (14)
where
  • $\Lambda=1$ (Adler-Moser) case: $l$ and $m$ are consecutive triangular numbers
    $l_{i}=i(2i+1),\quad m_{i}=i(2i-1).$ (15)
    In terms of standard notations for the Adler-Moser polynomials $P_{n}$, we have $p_{i}=P_{2i}$ and $q_{i}=P_{2i+1}$, and (9) writes as
    $P_{n}^{\prime\prime}P_{n-1}-2P_{n}^{\prime}P_{n-1}^{\prime}+P_{n}P_{n-1}^{\prime\prime}=0,\quad\deg(P_{n})=n(n+1)/2.$ (16)
  • $\Lambda=2$ case:
    $l_{i}=i(3i+2),\quad m_{i}=\frac{i(3i-1)}{2}.$ (17)
Now, consider bilinear relation (8) with $p=p_{i}$ and $q=q_{i}$ as a linear second-order ODE with solution $q_{i}$. Its second linearly independent solution $q=q_{i+1}$ equals
[…]
Similarly, considering $p_{i-1}$ and $p_{i}$ as linearly independent solutions, we get
[…]
By freedom in choosing linearly independent solutions of (8) one can also write analogs of (18) and (19) for decreasing indices
[…]
and one can generate $p_{i}$, and $q_{i}$ iteratively in either directions starting at some $i$.
Rewriting (18,19) or (20) in the differential form, we get the first-order differential recurrence relations
$\begin{array}[]{ll}q^{\prime}_{i+1}q_{i}-q_{i+1}q^{\prime}_{i}=\left(\frac{2l_{i}}{\Lambda}-2m_{i}+1\right)p_{i}^{2/\Lambda},\\ p^{\prime}_{i}p_{i-1}-p_{i}p^{\prime}_{i-1}=\left(2\Lambda m_{i}-2l_{i-1}+1\right)q_{i}^{2\Lambda}\end{array}.$ (21)
In the case $\Lambda=1$, the above recurrence relations are the three-term relations for the Adler-Moser polynomials $P_{n}$ with $q_{i}=P_{2i}$ and $p_{i}=P_{2i+1}$:
$P^{\prime}_{i+1}P_{i-1}-P^{\prime}_{i-1}P_{i+1}=(2i+1)P^{2}_{i}.$ (22)
Examples of several first Adler-Moser polynomials are
$P_{0}=1,\quad P_{1}=z,\quad P_{2}=z^{3}+t_{1},\quad P_{3}=z^{6}+5t_{1}z^{3}+t_{2}z-5t_{1}^{2},\quad\dots,$ (23)
where $t_{i}$ are free parameters related to the KdV “times” [1] (we set parameter related to shift of $z$ to zero).
In the $\Lambda=2$ case we get
$\begin{array}[]{ll}q^{\prime}_{i+1}q_{i}-q_{i+1}q^{\prime}_{i}=(3i+1)p_{i},\\ p^{\prime}_{i}p_{i-1}-p_{i}p^{\prime}_{i-1}=\left(6i-1\right)q_{i}^{4},\end{array}$ (24)
where we used (17). In difference from the Adler-Moser case, there are two distinct branches of polynomials both starting at $p_{0}=1$ and $q_{0}=1$. One branch with $i\geq 0$ going in the positive direction
$\begin{array}[]{ll}q_{0}=1&\quad p_{0}=1\\ q_{1}=z&\quad p_{1}=z^{5}+t_{1}\\ q_{2}=z^{5}+s_{2}z-4t_{1}&\begin{array}[]{r}\;\;p_{2}={z}^{16}+{\frac{44}{7}}s_{{2}}{z}^{12}-32t_{{1}}{z}^{11}+22{s_{{2}}}^{2}{z}^{8}-{\frac{2112}{7}}t_{{1}}s_{{2}}{z}^{7}\\ +1408{t_{{1}}}^{2}{z}^{6}+t_{{2}}{z}^{5}-44{s_{{2}}}^{3}{z}^{4}+352t_{{1}}{s_{{2}}}^{2}{z}^{3}\\ -1408s_{{2}}{t_{{1}}}^{2}{z}^{2}+2816{t_{{1}}}^{3}z+t_{{2}}t_{{1}}-{\frac{11}{5}}{s_{{2}}}^{4}\end{array}\\ \dots&\dots\end{array}$ (25)
etc, and another branch with $i\leq 0$ going in the negative direction
$\begin{array}[]{lll}p_{0}=1&&q_{0}=1\\ p_{-1}=z&&q_{-1}=z^{2}+s_{-1}\\ \begin{aligned} p_{-2}=z^{8}+\frac{28}{5}s_{-1}z^{6}+14s_{-1}^{2}z^{4}\\ +28s_{-1}^{3}z^{2}+t_{-2}z-7s_{-1}^{4}\end{aligned}&&\begin{aligned} q_{-2}=z^{7}+7s_{-1}z^{5}+35s_{-1}^{2}z^{3}+s_{-2}z^{2}\\ -35s_{-1}^{3}z+s_{-1}s_{-2}-\frac{5}{2}t_{-2}\end{aligned}\\ \dots&&\dots\end{array}$ (26)
etc, where $t_{i}$ and $s_{i}$ stand for arbitrary complex constants  [^2] Without loss of generality we omit a constant corresponding to a shift of $z$, so that $q_{1}=p_{-1}=z$. emerging in course of integrations in (18,19) or (20).
Summarizing, we emphasise again that, for $\Lambda=1$ and $\Lambda=2$, any no-common root solutions of the bi-linear equation (8) can be obtained by the recursive procedure described in this section. Therefore, the main sequences (23) or (26,25) determine all possible static patterns of a finite number of vortices with ${\cal Q}_{i}\in\{-1,1\}$ or ${\cal Q}_{i}\in\{-1,2\}$ respectively.
4 Darboux Transformations and Factorization Chains
The above results can be alternatively obtained with the help of factorization methods. The approach relying on these methods is more general than that used in the previous section and allows to obtain new stationary configurations. It involves theory of one-dimensional Schrodinger operator (“quantum Hamiltonian”) with potential $u(z)$:
$H=-\partial_{z}^{2}+u(z).$
Let us now take the operator
$H-\eta=-\partial_{x}^{2}+u(z)-\eta,$ (27)
and factorize into the product of two first-order operators
$H-\eta=A^{*}A,$ (28)
where [^3] Note that we do not use the special symbol $\circ$ for composition of operators and instead e.g. $A^{*}\circ A$ or $\varkappa\circ\partial_{z}\circ\varkappa^{-1}$ etc we simply write $A^{*}A$ or $\varkappa\partial_{z}\varkappa^{-1}$ etc.
$A^{*}=-\varkappa^{-1}\partial_{z}\varkappa=-\partial_{z}-\varkappa^{\prime}/\varkappa,\quad A=\varkappa\partial_{z}\varkappa^{-1}=\partial_{z}-\varkappa^{\prime}/\varkappa.$ (29)
The function $\varkappa=\varkappa(z)$ is an arbitrary element of kernel of $H-\eta$, i.e. $\varkappa$ is any eigenfunction of $H$ corresponding to the eigenvalue $\eta$
$H\varkappa=\eta\varkappa.$
It trivially follows from (28) that for any $\lambda$
$A(H-\lambda)=(\hat{H}-\lambda)A,$ (30)
where the new Schrodinger operator $\hat{H}$ is obtained from $H$ by permutation of factors $A^{*}$ and $A$ in (28):
$\hat{H}=AA^{*}+\eta=-\partial_{x}^{2}+\hat{u},\quad\hat{u}=u-2(\log\varkappa)^{\prime\prime}.$ (31)
Now, let us take an eigenfunction $\psi$ corresponding to the eigenvalue $\lambda$ of the old Schrodinger operator, i.e. $H\psi=\lambda\psi$. Then, from (30) it easy to see that the following transform of $\psi$:
$\hat{\psi}=A\psi=\psi^{\prime}-\frac{\varkappa^{\prime}}{\varkappa}\psi,$ (32)
is an eigenfunction of the new operator $\hat{H}$ corresponding to the same eigenvalue $\lambda$, i.e.
$\hat{H}\hat{\psi}=\lambda\hat{\psi}.$
In other words, we can obtain a new operator $\hat{H}$ and its eigenfunctions $\hat{\psi}(z;\lambda)$ from the old operator $H$ and its eigenfunctions $\psi(z;\lambda)$ corresponding to the same eigenvalues $\lambda$. New operator and eigenfunctions are obtained with the help of an eigenfunction $\varkappa=\psi(z;\eta)$ of the old operator $H$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (31,32) is called the Darboux Transformation.
When $\lambda\not=\eta$ the transform (32) of the two dimensional kernel of $H-\lambda$ is also two-dimensional. However, when $\lambda=\eta$, the transformation (32) annihilates the eigenfunction $\varkappa$ of $H$, i.e. $A\varkappa=0$: It maps the two-dimensional kernel $C_{1}\varkappa+C_{2}\varkappa\int\frac{dz}{\varkappa^{2}}$ of $H-\eta$ to the one dimensional sub-space
$A\left[C_{1}\varkappa+C_{2}\varkappa\int\frac{dz}{\varkappa^{2}}\right]=\frac{C_{2}}{\varkappa}.$
From the above, we see that function $1/\varkappa$ is an eigenfunction of the new Schrodinger operator corresponding to the eigenvalue $\eta$:
$\hat{H}\varkappa^{-1}=\eta\varkappa^{-1}.$
Then, the general solution $\hat{\varkappa}$ of $\hat{H}\hat{\varkappa}=\eta\hat{\varkappa}$ is
[…]
where the primitive of $\varkappa^{2}$ includes an arbitrary constant of integration. This extends the Darboux transformation of an eigenfunction (32) to the case $\lambda=\eta$. Transformations (33) are usually considered for $\eta=0$ and are called the Darboux transformations at the zero energy level.
Darboux transformations can be iterated. Sequence of iterated transformations is presented by the factorization chain:
$H_{0}=A_{0}^{*}A_{0}+\eta_{1}\to H_{1}=A_{0}A_{0}^{*}+\eta_{1}=A_{1}^{*}A_{1}+\eta_{2}\to H_{2}=A_{1}A_{1}^{*}+\eta_{2}=A_{2}^{*}A_{2}+\eta_{3}\to\dots.$ (34)
When all $\eta_{i}=0$, (34) becomes the chain of transformations at the zero energy level [^4] It would be more correctly to call it the chain of the Darboux transformations associated with zero energy level, rather than at the zero energy level, since eigenfunctions at all energy levels are transformed: Those with zero eigenvalue transform according to (33) and the rest according to (32).
$H_{0}=A_{0}^{*}A_{0}\to H_{1}=A_{0}A_{0}^{*}=A_{1}^{*}A_{1}\to H_{2}=A_{1}A_{1}^{*}=A_{2}^{*}A_{2}\to\dots,$ (35)
where
$A_{i}=\varkappa_{i}\partial_{z}\varkappa_{i}^{-1},\quad A_{i}^{*}=-\varkappa_{i}^{-1}\partial_{z}\varkappa_{i}.$
According to (33) and (31)
$\varkappa_{i+1}=\frac{C_{i}}{\varkappa_{i}}\int\varkappa_{i}^{2}dz,,\quad u_{i+1}=u_{i}-2(\log\varkappa_{i})^{\prime\prime},\quad H_{i}\varkappa_{i}=0,$ (36)
where $C_{i}$ are arbitrary constants. Using (36), we can construct iteratively the chain (35). Note, that at each step of the chain a free parameter, which is a constant of integration, appears in (36).
It trivially follows from (34) that $H_{0}$ and $H_{n}$ are related by the intertwining operator $T_{n}$,
$H_{n}T_{n}=T_{n}H_{0},\quad T_{n}=A_{n-1}A_{n-2}\cdots A_{0}.$ (37)
The function $T_{n}\psi(x;\lambda)$, where $\psi$ is an eigenfunction of $H_{0}$, ia an eigenfunction of $H_{n}$. Due to the Crum therem [10], the intertwining operator as well as potential $u_{n}$ can be explicitly written down in terms of eigenfunctions of the initial hamiltonian $H_{0}$. In more detail, the theorem states that the composition of $n$ Darboux transformations, corresponding to the factorization chain (34), transforms the eigenfunction $\psi$, corresponding to the eigenvalue $\lambda$ of $H_{0}=-\partial_{x}^{2}+u_{0}$ to the eigenfunction $\hat{\psi}$ of $H_{n}=-\partial_{x}^{2}+u_{n}$, corresponding to the same eigenvalue, as the ratio of two wronskians
$\hat{\psi}=\frac{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n},\psi]}{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n}]}$ (38)
while the potential is transformed in the following way
$u_{n}=u_{0}-2\left(\log{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n}]\right)^{\prime\prime}.$ (39)
Here $\psi_{i}$ are $n$ eigenfunctions of the initial Schrodinger operator $H_{0}$ corresponding to the distinct eigenvalues $\eta_{1},\eta_{2},\dots,\eta_{n}$, i.e. $H_{0}\psi_{i}=\eta_{i}\psi_{i}$, $1\leq i\leq n$, called “seed functions”, and ${\cal W}$ stands for the Wronskian determinant. From (38), we get the Wronskian representation of the intertwining operator [^5] Simplest proof of the Crum theorem uses the fact that, according to (32,37), $T_{n}$ annihilates $n$ seed functions $\psi_{i}$, $i=1,\dots,n$, i.e. kernel of $T_{n}$ is spanned by these seed functions. Then, since $T_{n}=\partial_{z}^{n}+\dots$, we immediately come to (40).
$T_{n}[\,\boldsymbol{\cdot}\,]=\frac{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n},\,\boldsymbol{\cdot}\,]}{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n}]}.$ (40)
In the particular case of confluent eigenvalues, when the spacing between $\eta_{i}$ vanishes, i.e. all $\eta_{i}$ tend to one value $\eta$ and we have the chain (35), the Crum theorem degenerates to
$\varkappa_{n}=\frac{{\cal W}[\varkappa,\partial_{\eta}\varkappa,\dots,\partial_{\eta}^{n-1}\varkappa,\partial^{n}_{\eta}\varkappa]}{{\cal W}[\varkappa,\partial_{\eta}\varkappa,\dots,\partial_{\eta}^{n-1}\varkappa]},\quad u_{n}=u_{0}-2\left(\log{\cal W}[\varkappa,\partial_{\eta}\varkappa,\dots,\partial_{\eta}^{n-1}\varkappa]\right)^{\prime\prime},$ (41)
where $\varkappa=\varkappa(z;\eta)$ is an eigenfunction of $H_{0}$ corresponding to the eigenvalue $\eta$, and $\varkappa_{n}$ is an eigenfunction of $H_{n}$ with the same eigenvalue. Here, the intertwining operator degenerates to
$T_{n}[\,\boldsymbol{\cdot}\,]=\frac{{\cal W}[\varkappa,\partial_{\eta}\varkappa,\dots,\partial_{\eta}^{n-1}\varkappa,\,\boldsymbol{\cdot}\,]}{{\cal W}[\varkappa,\partial_{\eta}\varkappa,\dots,\partial_{\eta}^{n-1}\varkappa]}.$
Eq. (41) is useful for representing the Adler-Moser polynomials in the Wronskian form. Indeed, let us take derivatives of the Schrodinger equation $H_{0}\varkappa(z,\eta)=\eta\varkappa(z,\eta)$, $H_{0}=-\partial_{z}^{2}+u_{0}(z)$ wrt $\eta$ and then set $\eta=0$. In this way, we get the recurrence relation for derivatives $\partial^{n}_{\eta}\varkappa$ at zero energy level:
$H_{0}\partial^{n}_{\eta}\varkappa|_{\eta=0}=n\partial^{n-1}_{\eta}\varkappa|_{\eta=0},\quad H_{0}\varkappa|_{\eta=0}=0.$
Thus, (41) can be rewritten as
$\varkappa_{n}=\frac{\tau_{n}}{\tau_{n-1}},\quad\tau_{n}={\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{n}],\quad u_{n}=u_{0}-2(\log\tau_{n-1})^{\prime\prime},\quad\tau_{0}=1,$ (42)
where the $\eta$-independent functions ${\cal X}_{n}(z)$ are defined by the recurrence relation
$H_{0}{\cal X}_{n}=c_{n}{\cal X}_{n-1},\quad H_{0}{\cal X}_{1}=0,\quad H_{0}=-\partial_{z}^{2}+u_{0}(z).$ (43)
In eq. (43), $c_{n}$ are arbitrary constants. At each iteration in (43) two integration constants appear, but only one of them is essential. The intertwining operator writes in terms of ${\cal X}_{i}$ as
$T_{n}[\,\boldsymbol{\cdot}\,]=\frac{{\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{n-1},\,\boldsymbol{\cdot}\,]}{{\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{n-1}]}.$ (44)
When we start the factorization chain (35) with the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$ (i.e. with $u_{0}=0$), from (43), we get
${\cal X}^{\prime\prime}_{n+1}(z)=(2n+1){\cal X}_{n}(z),\quad{\cal X}_{1}=z.$ (45)
Here [^6] In (45) we set $c_{n}=2n+1$. For such choice of normalization constants, $\tau_{n}$ is monic in $z$., the function ${\cal X}_{n}$ is a polynomial of degree $12n-$ 1 in $z$. Therefore $\tau_{n}$ in (42) is a polynomial of degree $n(n+1)/2$. Substituting $\varkappa_{n}=\tau_{n}/\tau_{n-1}$ (see(42)) into (36), we obtain recurrence relation for $\tau_{n}$ which coincides with that (22) for the Adler-Moser polynomials $P_{n}$. Taking into account that degrees of $\tau_{n}$ and $P_{n}$ are the same, we come to the conclusion that $\tau_{n}=P_{n}$. Thus, we get the Wronskian representation for the Adler-Moser polynomials
$P_{n}={\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{n}],$
where ${\cal X}_{n}$ are defined by (45).
5 Darboux Transformation and Equilibrium Configurations
As we mentioned above, sequences of the equilibrium configurations can be produced with help of Darboux transformations. To construct them, we need to introduce the factorizable function corresponding to a static configuration (6)
$\psi(z)=C\prod_{i=1}^{N}(z-z_{i})^{{\cal Q}_{i}},$ (46)
where $C$ is a non-zero constant and $Q_{i}$ are not necessarily integers.
Now, we will show that if the Darboux transform $\hat{\psi}$ of $\psi$ has a similar form, i.e.
$\hat{\psi}(z)=\hat{C}\prod_{i=1}^{\hat{N}}(z-\hat{z}_{i})^{\hat{\cal Q}_{i}},$ (47)
then $\hat{\psi}$ also corresponds to a static configuration. For this we introduce the Schrodinger operator $H$ whose kernel contains $\psi$, i.e.
$H\psi=(-\partial_{z}^{2}+u)\psi=0.$
Then, using (46), we obtain
$u=\frac{\psi^{\prime\prime}}{\psi}=(\log\psi)^{\prime\prime}+((\log\psi)^{\prime})^{2}=-\sum_{i}\frac{{\cal Q}_{i}}{(z-z_{i})^{2}}+\sum_{i,j}\frac{{\cal Q}_{i}{\cal Q}_{j}}{(z-z_{i})(z-z_{j})}.$
Applying the partial fraction decompositions in the last sum we get
$u=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}}+2\sum_{i}\frac{{\cal Q}_{i}}{(z-z_{i})}\sum_{j\not=i}\frac{{\cal Q}_{j}}{(z_{i}-z_{j})}.$
Due to the equilibrium condition (6) the last term in the above equation vanishes and potential is a sum of the second order poles
$u=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}},$ (48)
i.e. the equilibrium condition is equivalent to absence of the first-order poles in the potential.
Now, to prove that the Darboux transform $\hat{\psi}$ of $\psi$ of the form (47) also correspond to an equilibrium configuration, i.e.
$\sum_{j=1,j\not=i}^{\hat{N}}\frac{\hat{\cal Q}_{j}}{\hat{z}_{i}-\hat{z}_{j}}=0,\quad i=1\dots\hat{N},$
we have to show that potential $\hat{u}$ of the Darboux transformed Schrodinger operator $\hat{H}$ is also a sum of the second order poles. Indeed, the Darboux transformed potential equals (see (33))
$\hat{u}=u-2(\log\psi)^{\prime\prime},$
and from (46) and (48) we obtain
$\hat{u}=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}+1)}{(z-z_{i})^{2}},$ (49)
i.e. $\hat{u}$ is a sum of second order poles and therefore $\hat{\psi}$ of the form (47) corresponds to a static configuration.
Note, that since
$\hat{u}=\sum_{i}\frac{{\cal\hat{Q}}_{i}({\cal\hat{Q}}_{i}-1)}{(z-\hat{z}_{i})^{2}},$
from (49), we see that the Darboux transformation acts on charges either of two ways
$\hat{\cal Q}=-{\cal Q},\quad\hat{\cal Q}={\cal Q}+1,$ (50)
i.e. it either (i) inverts a charge or (ii) increments a charge by 1. In particular, it can create new charges with ${\cal Q}_{i}=1$ and annihilate charges with ${\cal Q}_{i}=-1$. In the case of the Adler-Moser polynomials in generic configurations, it annihilates negative charges, inverts positive charges, and creates positive charges at new positions.
6 Configurations Related to Bi-spectral Systems
In section 3 we showed that the static configurations of the system consisting of charges of values ±$\pm 1$ (i.e. when $\Lambda=1$) correspond to zeros of the Adler-Moser polynomials $P_{n}$ and $P_{n+1}$. In section 4 these polynomials were obtained with the help of sequence of the Darboux transformations (35) at zero energy level. Corresponding sequence of $H_{i}$ is started from the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. This sequence is known to constitute the “odd” family of bi-spectral operators [^7] The differential operators $H$ in $z$ is bi-spectral if there exists a differential operator $B$ in $k$ and a common eigenfunction, such that $H\psi(z,k)=f(k)\psi(z,k)$, $B\psi(z,k)=g(z)\psi(z,k).$ For the even family of Schrodinger operators, the common eigenspace of $H$ and $B$ is two-dimensional. It is one-dimensional in the odd case. [17]. There exists another sequence of the Darboux transforms consisting of “even” family, of bi-spectral operators [17]. It starts with
$H_{0}=-\partial_{z}^{2}-\frac{1}{4z^{2}}.$ (51)
In both “even” and “odd” cases we present elements of the sequence of Darboux-transforms of kernels of $H_{i}$
$\psi_{0}\to\psi_{1}\to\psi_{2}\to\dots,\quad H_{i}\psi_{i}=0$
as the ratios
$\psi_{i}=P_{i+1}/P_{i},\quad P_{0}=1.$ (52)
According to section 4 (see eq. (33)) the Darboux transform $\psi_{i+1}$ of $\psi_{i}$ is
[…]
From this and (52) we get the differential recurrent relation for $P_{i}$
$P^{\prime}_{i+1}P_{i-1}-P^{\prime}_{i-1}P_{i+1}=C_{i}P^{2}_{i}.$ (54)
In the “odd”, i.e. the Adler-Moser case $P_{0}=1$, $P_{1}=z$ (see eq. (22)). In the “even” case $H_{0}[z^{1/2}]=0$ (i.e. $\psi_{0}=z^{1/2}$ see (51)) and, according to (52)
$P_{0}=1,\quad P_{1}=z^{1/2}.$ (55)
Note that, for convenience, we set the normalization constants $C_{i}$ in (54) to values at which $P_{i}(z)$ are monic [^8] In the even case, this differs from the original normalization of [17], where same $C_{i}=2i+1$ were chosen for both families.. For “even” family, from (54,55) we get $P_{2}=z^{2}+t_{1}$. So, using (52), we obtain the sequence of length 2 depending on one parameter $t_{1}$:
$\psi_{0}=z^{1/2},\quad\psi_{1}=\frac{z^{2}+t_{1}}{z^{1/2}}.$ (56)
Another application of the recurrence relation (54) with $t_{1}\not=0$ would produce a logarithmic term, so to obtain $\psi_{2}$ and $\psi_{3}$ of the form (46) we restrict $t_{1}$ to be zero and pick a new integration constant $t_{2}$:
$P_{3}=z^{9/2}+t_{2}z^{1/2},\quad P_{4}=z^{8}+6t_{2}z^{4}+t_{3}z^{2}-3t_{2}^{2}.$ (57)
Now, we get the sequence of length 4:
$\psi_{0}=z^{1/2},\quad\psi_{1}=z^{3/2},\quad\psi_{2}=\frac{z^{4}+t_{2}}{z^{3/2}},\quad\psi_{3}=\frac{z^{8}+6t_{2}z^{4}+t_{3}z^{2}-3t_{2}^{2}}{z^{9/2}+t_{2}z^{1/2}}$
depending on two parameters $t_{2}$,$t_{3}$.
Next, to eliminate logarithmic terms in $P_{5}$ and $P_{6}$, one has to set both $t_{1}=t_{2}=0$ and we get a sequence $\psi_{0},\psi_{1},\psi_{2},\psi_{3},\psi_{4},\psi_{5}$ (of length 6) depending on three parameters $t_{3}$, $t_{4}$ and $t_{5}$. And so forth.
Sequences of the even family have finite lengths (i.e. they terminate) due to the presence of charge of the third specie at the origin: At the beginning of sequence, charge $1/2$ is placed at $z=0$. Then, according to (50), the following transformations are produced with this charge: It is incremented by $1$ during each of the $m-1$ first Darboux transformations. It is inverted by the $m$th transformation and becomes equal to $-m+1/2$. Then, again, it is incremented during the next $m-1$ steps, until it reaches value $-1/2$. This charge would create a logarithmic singularity in (53) at the next step, so sequence terminates.
Closing this section we note that similarly to the odd family, the wronskian representation of $\psi_{i}$ exists also for the even family. It will be given in section 11 (eqs. (88) and (89)).
7 KWCC Transformation and Terminating Configurations
Let us now return to the sequence of transformations of rational functions (whose primitives are also rational functions) corresponding to  [^9] Or corresponding to the chain of solutions $\dots\to p_{i-1},q_{i}\to p_{i},q_{i}\to p_{i},q_{i+1}\to p_{i+1},q_{i+1}\to\dots$ of (8). iterations (18,19)
$\dots\to\frac{q_{i}^{2\Lambda}}{p_{i-1}^{2}}\to\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}\to\frac{q_{i+1}^{2\Lambda}}{p_{i}^{2}}\to\frac{p_{i+1}^{2/\Lambda}}{q_{i+1}^{2}}\to\dots.$ (58)
Denoting function $\frac{q_{i}^{2\Lambda}}{p_{i-1}^{2}}$ from the above chain as $\psi^{2}$ and the next function $\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}$ as $\tilde{\psi}^{2}$, from (19) we get the relationships between them
[…]
where, for this step of the chain, $\gamma=1/\Lambda$ and $C$ stands for a non-zero constant.
Next transformation in the sequence (58) relates $\psi^{2}=\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}$ and $\tilde{\psi}^{2}=\frac{q_{i+1}^{2\Lambda}}{p_{i}^{2}}$. Using (18) we also arrive to transformation (59), this time with $\gamma=\Lambda$.
Therefore, the chain (58) can be constructed by the iterative applications of following transformations
[…]
where
$\gamma_{i+1}=1/\gamma_{i},$ (61)
and
$\gamma_{i}\in\{1/2,1,2\}.$ (62)
The transformation (59) was presented by Krishnamurthy, Wheeler, Crowdy and Constantin (KWCC) in [22].
The KWCC transformation (59) is a composition of two mappings
  1. (i)
[…]
  2. (ii)
    Exponentiation $\tilde{\psi}=\hat{\psi}^{\gamma}$.
[…]
From the above and (50) it follows that KWCC transformation (59) acts on charges either of two ways
$\tilde{\cal Q}=-\gamma{\cal Q},\quad\tilde{\cal Q}=\gamma\left({\cal Q}+1\right).$ (63)
Obviously, the KWCC transformations (60,61,62) generate all classes of configurations considered above, i.e. those related to the odd and even bi-spectral families as well as to the $\Lambda=2$ case.
To rewrite the sequence of transforms (60), (61) in terms of polynomials, we present $\psi_{i}$ as
$\psi_{i}=\frac{\tau_{i}^{\gamma_{i}}}{\tau_{i-1}},$
where $\tau_{i}$ is a polynomial of an integer or half-integer degree $d_{i}$. Here, “polynomial of half-integer degree” in $z$ stands for a polynomial in $z$ times $z^{1/2}$. Then from (60) and (61) we get
[…]
where
$d_{i+1}=2\gamma_{i}d_{i}-d_{i-1}+1,\quad\gamma_{i+1}=1/\gamma_{i},\quad d_{0}=0.$ (64)
In the differential form these relations write as
$\tau_{i+1}^{\prime}\tau_{i-1}-\tau_{i+1}\tau_{i-1}^{\prime}=(2\gamma_{i}d_{i}-2d_{i-1}+1)\tau_{i}^{2\gamma_{i}},\quad\tau_{i}(z)=z^{d_{i}}+\dots,\quad\tau_{0}=1.$ (65)
These are actually relations (21), but now $d_{i}$ are not necessarily degrees (15) or (17) of the main sequences. This is because (64, 65) are derived from the KWCC chain, rather than from the bi-linear equation (8). Because of this fact, any polynomial solution of (65) that starts from a static configuration (e.g. from a single vortex) also corresponds to a sequence of static configurations.
Apart from infinite main sequences of the $\Lambda=1$ and $\Lambda=2$ cases, recurrence relations (64,65) allow to find sequences of configurations of finite lengths, i.e. terminating configurations. The “even” bi-spectral family of section 6, i.e. solutions of (64,65) corresponding to $\gamma_{i}=1$, $d_{1}=1/2$, is a family of terminating configurations for $\Lambda=1$.
As an example of family of terminating configurations for $\Lambda=2$, we can take the case $\gamma_{0}=2$, $d_{1}=2$, i.e. $\tau_{1}=z^{2}$. Here are several first terminating sequences of $\tau_{i}$:
$\begin{array}[]{c}1,z^{2},z^{3}+t_{1}\\ 1,z^{2},z^{3},z^{2}(z^{9}+t_{2}),z^{9}+t_{3}z^{3}-2t_{2}\\ 1,z^{2},z^{3},z^{11},z^{3}(z^{6}+t_{3}),z^{2}(z^{24}+\frac{20}{3}z^{18}t_{3}+30z^{12}t_{3}^{2}+z^{9}t_{4}-20z^{6}t_{3}^{3}-\frac{5}{3}t_{3}^{4}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad z^{18}+15z^{12}t_{3}+z^{9}t_{5}-45z^{6}t_{3}^{2}+\frac{2t_{3}t_{5}-3t_{4}}{2}z^{3}+5t_{3}^{3}\end{array}$
Similarly to the even bi-spectral family, here, a charge of the third specie undergoes transformations (63) at $z=0$ (this time with $\gamma\in\{1/2,2\}$) until the logarithmic singularity is encountered.
One can compute terminating sequences by solving (65) for $\tau_{i+1}$ as for a monic polynomial with unknown coefficients. Compatibility of resulting system of algebraic equations (which is linear in coefficients of $\tau_{i+1}$) imposes restrictions on free parameters of $\tau_{i}$. For sequences related to the even bi-spectral family there exists a wronskian representation (see section 11).
We recall that the main sequences exhaust all possible no-common root polynomial solutions of (8) as well as the collapse solutions for $\Lambda=1$ and $\Lambda=2$ cases. Solutions of (65) whose degrees are not in the main sequences do not satisfy bi-linear equations (8). They satisfy its generalization.
In more detail: Let us first consider sequence of the Darboux transforms at zero energy level that starts from $H_{0}=-\partial_{z}^{2}+u_{0}$, with non-zero $u_{0}\not=0$. At the $n$th step of the chain we have
$u_{n}=u_{0}-2(\log\tau_{n})^{\prime\prime},\quad\varkappa_{n}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{0}=1,\quad H_{n}\varkappa_{n}=0,$
Therefore
$H_{n}\varkappa_{n}=\left(-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}\right)\left[\frac{\tau_{n+1}}{\tau_{n}}\right]=0,$
or equivalently
$\tau_{n+1}^{\prime\prime}\tau_{n}-2\tau_{n+1}^{\prime}\tau_{n}^{\prime}+\tau_{n+1}\tau_{n}^{\prime\prime}-u_{0}\tau_{n}\tau_{n+1}=0.$ (66)
This is a generalization of the Tkachenko equation (9) that includes cases of non-integer $d_{i}$. For the even bi-spectral family, the initial potential equals to $u_{0}=a(a-1)/z^{2}$, where $a$ is a half-integer. Then, substituting $\tau_{n}=p$, $\tau_{n+1}=z^{a}q$ into (66), we get
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}+\frac{2a}{z}(q^{\prime}p-p^{\prime}q)=0.$
In fact, this generalization is a particular case of the tri-linear equation for three species of vortices with ${\cal Q}_{i}\in\{-1,1,a\}$, where a single vortex of the third specie of strength $a$ is placed at the origin $z=0$ [15, 33]. In general, for multiple species we have the multi-linear equation (see e.g. [25, 32])
$\sum_{i}\Lambda_{i}^{2}\frac{{\cal P}_{i}^{\prime\prime}}{{\cal P}_{i}}+2\sum_{i<j}\Lambda_{i}\Lambda_{j}\frac{{\cal P}_{i}^{\prime}}{{\cal P}_{i}}\frac{{\cal P}_{j}^{\prime}}{{\cal P}_{j}}=0,$
where sums run over all species, $\Lambda_{i}$ is a strength of the $i$th specie and ${\cal P}_{i}$ is polynomial whose roots are coordinates of vortices of this specie. In the case of configurations related to the even bi-spectral family ${\cal P}_{1}=p$, ${\cal P}_{2}=q$, ${\cal P}_{3}=z$. In similar situation, when ${\cal Q}_{i}\in\{-1,\Lambda,\Lambda a\}$, we have the following generalization of (8)
$p^{\prime\prime}q-2\Lambda p^{\prime}q^{\prime}+\Lambda^{2}pq^{\prime\prime}+\frac{2a\Lambda}{z}(\Lambda q^{\prime}p-p^{\prime}q)=0.$
Exploring this equation, O’Neil and Cox-Steib [33] presented the families of terminating configurations.
It is worth mentioning that all families of equilibrium patterns obtained above are related to the following chain of transformations that starts from the free Schrodinger operator
$A_{-1}A_{-1}^{*}\xrightarrow{\gamma_{0}}A_{0}^{*}A_{0}\to A_{0}A_{0}^{*}\xrightarrow{\gamma_{1}}A_{1}^{*}A_{1}\to\dots,\quad\gamma_{i}\in\{1/2,1,2\},$
where $A_{-1}=-A^{*}_{-1}=\partial_{z}$, and $\xrightarrow{\gamma_{i}}$ stand for exponentiations of kernels. For example, for the even bi-spectral family $\gamma_{0}=1/2$ and $\gamma_{i}=1$, $i>0$. Chains of this type are generalizations of (35).
8 Translating Configurations and Baker-Akhieser Functions
Translating configuration of vortices are solutions of (8). To generate them we consider the following eigenfunction
$\psi(k,z)=C\prod_{i=1}^{N}(z-z_{i}(k))^{{\cal Q}_{i}}e^{kz},$ (67)
of the Schrodinger operator
$H\psi=(-\partial_{z}^{2}+u)\psi=-k^{2}\psi.$
The condition that potential $u=\psi^{\prime\prime}/\psi-k^{2}$ does not have simple poles (i.e. it is of form (48)) is equivalent to the equilibrium conditions (8) for a translating configuration. This can be shown by arguments similar to that used for the static configurations in Section 5.
In the case of the factorization chain (35) starting with the free Schrodinger operator, there exists an eigenfunction, called the “rational Baker-Akhieser function” [^10] In theory of integrable hierarchies of non-linear PDEs, the Baker-Akhieser function is a solution of an infinite system of linear differential equations. Associated integrable hierarchy arises as a condition of compatibility of the system. Here we deal with the Baker-Akhieser function of the KdV hierarchy., which has the form (67). It is an eigenfunction of $H_{j}$ that is obtained by the action of an intertwining operator with rational coefficients on the eigenfunction $e^{kz}$ of the free Schrodinger operator, i.e. $\Psi=C(k)T_{j}[e^{kz}]$. Since the $z$-independent factor $C(k)$ does not play role in our considerations, we will use the eigenfunction
$\psi_{j}=T_{j}[e^{kz}]=\frac{q_{j}}{p_{j}}e^{kz},$ (68)
where $p_{j}$ and $q_{j}$ are polynomials in $z$. As we have shown in section 6, the corresponding potential $u_{j}$, obtained by $j$ Darboux transformations from $u_{0}=0$, is of the form (48). Since $u_{j}$ does not contain simple poles, eigenfunction (68) corresponds to a translating configuration. It will be shown in section 12 that the class of such configurations contains all possible translating patterns consisting of finite number of vortices with ${\cal Q}_{i}=\pm 1$.
A constructive way to write down $\psi_{j}$ is to use the Wronskian representation of $T_{j}$, given by (44, 45). Then, from (68) we get
$\psi_{j}=\frac{{\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j},e^{kz}\right]}{{\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j}\right]},\quad{\cal X}_{1}=z,\quad{\cal X}^{\prime\prime}_{j}=(2j+1){\cal X}_{j-1},$
and from (68) it follows that
$q_{j}=e^{-kz}{\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j},e^{kz}\right],\quad p_{j}={\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j}\right]=P_{j},$ (69)
where $P_{j}$ is the $j$th Adler-Moser polynomial. When $k\not=0$, degree of $q_{j}(z)$ equals to that of $p_{j}(z)$. In the electrostatic interpretation (see section 1), this means that the total charge has to be zero $\sum_{i}{\cal Q}_{i}=0$ in order the system to be at rest in the homogeneous electric field $k\not=0$. In the $k\to 0$ limit
$\psi_{j}(k=0,z)={\rm const}\frac{P_{j-1}}{P_{j}},$
and we get the static equilibrium with non-zero total charge. In other words, in the $k\to 0$ limit a part of positive charges moves to infinity.
9 Equilibria in Non-Homogenous Background Fields
Let us assume that the Schrodinger operator $H=-\partial_{z}^{2}+u(z)$ has the following “quasi-rational” eigendunction
$\psi(z)=C\prod_{i=1}^{N}(z-z_{i})^{{\cal Q}_{i}}e^{\Phi(z)},\quad H\psi=\lambda\psi,$ (70)
where $\Phi(z)$ is a polynomial. “Quasi-rational” here means that all $Q_{i}$ are integers. Requirement that potential $u=\frac{\psi^{\prime\prime}}{\psi}+\lambda$ does not have simple poles results in the system of algebraic equations
$\Phi^{\prime}(z_{i})+\sum_{i=1,i\not=j}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}=0,\quad i=1\dots N.$ (71)
Equations (71) are conditions of equilibrium of charges in an external harmonic electric field [^11] Such conditions were first presented by Stieltjes [36, 37] for roots of orthogonal polynomials. or equilibrium of vortices in a background flow, see section 1. Function $\Phi(z)$ is a potential of the external field/background flow. Provided (71) holds,
$u(z)=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}}+\Phi^{\prime\prime}(z)+\Phi^{\prime}(z)^{2}+2\sum_{i}{\cal Q}_{i}\frac{\Phi^{\prime}(z)-\Phi^{\prime}(z_{i})}{z-z_{i}}-\lambda.$
If there exists another quasi-rational eigenfunction of $H$ which is of the form similar to (70)
$\tilde{\psi}(z)=C\prod_{i=1}^{\tilde{N}}(z-\tilde{z}_{i})^{{\cal\tilde{Q}}_{i}}e^{\Phi(z)},\quad H\tilde{\psi}=\eta\tilde{\psi},$
it can be taken as a “seed” for the Darboux transformation. Then the transform $\hat{\psi}$ of $\psi$
$\hat{\psi}=\psi^{\prime}-\frac{\tilde{\psi}^{\prime}}{\tilde{\psi}}\psi,\quad\hat{H}\hat{\psi}=(-\partial_{z}^{2}+\hat{u})\hat{\psi}=\lambda\hat{\psi}$ (72)
will be also quasi-rational of the form similar to (70), i.e.
$\hat{\psi}(z)=C\prod_{i=1}^{\hat{N}}(z-\hat{z}_{i})^{{\cal\hat{Q}}_{i}}e^{\Phi(z)}.$
Since the transform of potential
$\hat{u}=u-2(\log\tilde{\psi})^{\prime\prime}$
does not have first-order poles, $\hat{\psi}$ corresponds to a new equilibrium configuration in the same external field. For a set of eigenfunctions of $H$ (“seed functions”) of the form
$\varkappa_{i}=R_{i}(z)e^{\Phi(z)},\quad H\varkappa_{i}=\lambda_{i}\psi_{i},$
where $R_{i}(z)$ are rational in $z$, we can apply several Darboux transformations resulting in various equilibrium configurations. According to the Darboux-Crum theorem (38,39), the results of compositions of $n$ transformations can be written in the Wronskian form
$\psi^{(n)}(z)=\frac{{\cal W}[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n-1}},\varkappa_{k_{n}}]}{{\cal W}[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n-1}}]}.$ (73)
For example, for harmonic oscillator $u=z^{2}/2$, $\Phi(z)=-z^{2}/2$ and $R_{n}$ are Hermite polynomials. In this case, transforms (73) correspond to various equilibria of charges in the external quadrupole field or vortex patterns in vicinity of the stagnation point of the flow encountering a solid boundary (for details and other examples, see e.g. [14, 25]). Note, that in difference from the systems considered in the previous sections, configurations corresponding to Hermite polynomials do not depend non-trivially on continuous parameters and are rather parametrised by sequences of distinct integers $k_{1}<k_{2}<\dots<k_{n}$.
10 Equilibria of Vortex Streets - Odd Family
A periodic vortex street is a vortex pattern consisting of an infinite number of vortices situated periodically along some direction. A vortex street can be equivalently considered as a configuration of $N$ vortices (or charges) on a cylinder. In what follows by the terms “vortex street”, “periodic configuration” or “periodic pattern” we mean configuration of a finite number of charges on cylinder. Without loss of generality we can set the translation period to $\pi$. Since $\sum_{n=-\infty}^{\infty}\frac{1}{z+\pi n}=\cot z$, the stationarity condition (5) for the translating (with a speed ${\rm i}\bar{k}/(2\pi)$) vortex street becomes
$k+\sum_{j=1,j\not=i}^{N}{\cal Q}_{j}\cot(z_{i}-z_{j})=0,\quad i=1\dots N,$ (74)
and for static periodic patterns (see (6)) we have
$\sum_{j=1,j\not=i}^{N}{\cal Q}_{j}\cot(z_{i}-z_{j})=0,\quad i=1\dots N.$ (75)
Similarly to vortex patterns on the plane that are related to the rational solutions of the KdV hierarchy or the odd bi-spectral family, the trigonometric soliton solutions of the hierarchy are related to the patterns on a cylinder [19, 21, 25]. The latter can be shown by methods similar to those used in sections 5, 8.
In more detail, we introduce the trigonometric-rational analog of (46) of the following form
$\psi(z)=C\prod_{i=1}^{N}\sin^{{\cal Q}_{i}}(z-z_{i}),$ (76)
where ${\cal Q}_{i}$, $z_{i}$ correspond to the periodic static configuration (75). The “trigonometric-rational” here means that all ${\cal Q}_{i}$ are integers. Trigonometric-rational are special cases of the trigonometric-factorizable functions, where ${\cal Q}_{i}$ are real. Function (76) is an eigenfunction of the Schrodinger operator
$H\psi=\lambda\psi,\quad H=-\partial_{z}^{2}+u(z)$
with the potential
$u(z)=\frac{\psi^{\prime\prime}(z)}{\psi(z)}+\lambda.$
Taking into account the identity
$\cot(z-z_{i})\cot(z-z_{j})=-1+\left(\cot(z-z_{i})-\cot(z-z_{j})\right)\cot(z_{i}-z_{j}),$
similarly to the rational case, we come to the conclusion that the static equilibrium condition is equivalent to absence of first-order poles in the potential and
$u=\sum_{i=1}^{N}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{\sin^{2}(z-z_{i})}+{\rm const}.$ (77)
A family of the trigonometric-rational configurations can be generated by the chain of Darboux transformations (34) with $\eta_{i}=k_{i}^{2}$, where $0<k_{1}<k_{2}<\dots<k_{i}<\dots$ is a sequence of positive integers. The chain starts from the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$ and it is constructed with the help of “seed functions” that are its periodic eigenfunctions
$\varkappa_{k_{i}}(z)=\sin(k_{i}z+\zeta_{i}).$ (78)
Then, according to the Crum theorem (38,39), at $n$th step of the chain we have
$\psi^{(n)}(z)=\frac{\tau_{n+1}(z)}{\tau_{n}(z)},\quad\tau_{n}={\cal W}\left[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n}}\right],\quad\tau_{0}=1.$ (79)
Obviously, $\tau_{n}(z)$ are trigonometric polynomials, $\psi^{(n)}(z)$ are trigonometric-rational functions and corresponding potentials
$u_{n}(z)=-2(\log\tau_{n}(z))^{\prime\prime}$ (80)
are of the form (77) and do not have simple poles. Thus, eigenfunctions (79) correspond to the static vortex street configurations. Note that potentials (80) are periodic $n$-soliton solutions of the KdV hierarchy and $\tau_{n}$ are its $n$-soliton $\tau$-functions (see e.g. [29] and references therein). We will demonstrate in section 12 that the set of these configurations actually contains all static vortex street patterns with ${\cal Q}_{i}=\pm 1$.
For translating configurations (see (74)) related to the soliton solutions we can apply metod similar to that used in section 8. By analogy with quasi-rational (67), we now consider the trigonometric quasi-rational eigenfunctions
$\psi(k,z)=C\prod_{i=1}^{N}\sin(z-z_{i}(k))^{{\cal Q}_{i}}e^{kz},\quad{\cal Q}_{i}\in\mathbb{Z}.$ (81)
According to the Darboux-Crum theorem (38), eigenfunctions, obtained by a sequence of the Darboux transformation from the formal eigenfunction $e^{kz}$ of the free Schrodinger operator, equal
$\psi^{(n)}(k,z)=\frac{{\cal W}\left[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n}},e^{kz}\right]}{{\cal W}\left[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n}}\right]},$ (82)
with $\varkappa_{k_{i}}$ given by (78). Obviously $\psi^{(n)}(k,z)$ are of the form (81) and therefore correspond to the translating vortex street configurations. Since each $\varkappa_{k_{i}}$ (78) depends on a free parameter, the translating and static configurations related to soliton solutions depend on $n$ free parameters as well as on the sequence of integers $0<k_{1}<k_{2}\dots<k_{n}$. Up to a $z$-independent common factor $C(k)$, (82) is the trigonometric Baker-Akhieser function of the KdV hierarchy [^12] It is also the “angular part” of the Baker-Akhieser function of 2d quantum algebraically integrable systems, see e.g. [5],[12] and references therein..
The class of configurations determined by (82,78) contains all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in section 12.
11 Equilibria of Vortex Streets - Even Family
In the previous section we considered vortex street configurations related to the trigonometric analogs of the odd family of bi-spectral potentials or soliton solutions of the KdV hierarchy. Here we consider trigonometric analogs of the even family (see section 6), i.e. examples of periodic configurations where not all ${\cal Q}_{i}$ are integers. For this, let us start our chain of Darboux transformations with the trigonometric Pochl-Teller potential
$u_{0}=\frac{a(a-1)}{\sin^{2}z}+\frac{b(b-1)}{\cos^{2}z}$ (83)
with arbitrary real $a$ and $b$. The simplest configuration corresponds to the eigenfunction $\sin^{a}z\cos^{b}z$ and consists of two static vortices of the strengths ${\cal Q}_{1}=a$, ${\cal Q}_{2}=b$ situated at the diametrically opposite points of the $\pi$-periodic cylinder, i.e. at $z_{1}=0$, $z_{2}=\pi/2$.
Note that when $a=b$, or $a=1-b$, the potential (83) can be rewritten as $u_{0}=4a(a-1)/\sin^{2}2z$. By re-scaling $2z\to z$ in the Schrodinger operator, we again come to the special $b=0$ case of the $\pi$-periodic potential (83).
For integers $a$ and $b$, (83) is a soliton solution of KdV at specific “times” (corresponding parameters $k_{i}$ and $\zeta_{i}$ in (79, 78) can be found e.g. in refs [6] or [18]) [^13] In general, sequence of $n$ Darboux transformations of the chain (34), where $A_{i}=\partial_{z}-(a+i)\cot z+(b+i)\tan z$, $i=0,1,\dots$., results in the shift of parameters $a\to a+n$, $b\to b+n$ in the potential (83).. In this case, $u_{0}$ and sequences of its Darboux transforms belong to the odd family considered in the previous section.
Due to invariance of the potential (83) wrt involutions $a\to 1-a$, $b\to 1-b$, when $a\pm b$ is not integer (generic $a$ and $b$), the Schrodinger operator $H_{0}=-\partial_{z}^{2}+u_{0}$ has four distinct types of factorizable eigenfunctions $\varkappa_{i}(z)$, $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ of the form
$\begin{array}[]{ll}{}_{2}F_{1}(-i,i+a+b,1/2+a;\sin^{2}z)\sin^{a}z\cos^{b}z,&\lambda_{i}=(2i+a+b)^{2},\\ {}_{2}F_{1}(-i,i+1-a+b,3/2-a\;\sin^{2}z)\sin^{1-a}z\cos^{b}z,&\lambda_{i}=(2i+1-a+b)^{2},\\ {}_{2}F_{1}(-i,i+a+1-b,1/2+a;\sin^{2}z)\sin^{a}z\cos^{1-b}z,&\lambda_{i}=(2i+1+a-b)^{2},\\ {}_{2}F_{1}(-i,i+2-a-b,3/2-a;\sin^{2}z)\sin^{1-a}z\cos^{1-b}z,&\lambda_{i}=(2i+2-a-b)^{2},\end{array}$ (84)
where $i=0,1,2,\dots$ and ${}_{2}F_{1}$ stands for the Gauss hypergeometric function. The first parameter of the ${}_{2}F_{1}$ in (84) is a non-positive integer $-i$, and the hypergeometric factors in (84) are polynomials in $\sin^{2}z$. These factorizable eigenfunctions are trigonometric-rational up to a factor $\sin^{\pm a}z\cos^{\pm b}z$. For generic $a$ and $b$ all eigenvalues $\lambda_{i}$ are distinct and all eigenfunctions in (84) are linearly independent.
From (84), with the help of the Wronskian formula (79), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and therefore correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in difference from the soliton-related configurations (79,78), configurations generated by (84) do not depend on free parameters others than $a$ and $b$. Similarly to systems considered in section 9, configurations with generic $a$ and $b$ are parametrized only by sequences of distinct integers $0<k_{1}<k_{2}<\dots<k_{n}$.
Even family of periodic solutions emerges when $a$ and $b$ become half-integers, such that
$a=l+1/2,\quad b=m+1/2,\quad l,m\in\mathbb{N}.$
In this case, eigenvalues $\lambda_{i}$ in (84) overlap, and some of them become degenerate, with two-dimensional eigenspaces [^14] Degeneration occurs also for integer $a$ and $b$. There, number of degenerate states, as we saw in the previous section, is infinite.. Now, there are two types (degenerate and non-degenerate) of eigenfunctions $\varkappa_{i}(z)$ :
$\begin{array}[]{lll}P^{(-l,-m)}_{i}(\cos 2z,\zeta_{i})\sin^{1/2-l}z\cos^{1/2-m}z,&\lambda_{i}=(2i+1-l-m)^{2},&\frac{l+m}{2}\leq i<l+m\\ P^{(l,m)}_{i}(\cos 2z)\sin^{l+1/2}z\cos^{m+1/2}z,&\lambda_{i}=(2i+1+l+m)^{2},&i\geq 0\end{array},$ (85)
which are both trigonometric-rational up to the common factor $(\sin z\cos z)^{1/2}$. The first type, $P^{(-l,-m)}_{i}(x,\zeta)$ is a polynomial in $x$ which depending also on a free parameter $\zeta$. Such polynomials are called para-Jacobi Polynomials [11], [8]. Solutions of the second type involve ordinary Jacobi polynomials $P^{(l,m)}_{i}$ and are non-degenerate. In difference from the latter, the para-Jacobi polynomials are not orthogonal.
Thus, the even family of the static vortex street configurations parametrized by sequences of continuous free parameters as well as sequences of integers can be constructed with the help of (79) and (85) for half-integer $a$, $b$.
The simplest non-trivial example of a member of the even family corresponds to $l=m=1$. Here $a=b=3/2$, $P^{(-1,-1)}_{1}(\cos 2z,\zeta_{1})=\cos 2z+1+\zeta_{1}$ and from (85) we get
$\psi_{1}=\varkappa_{1}=\frac{\cos 2z+1+\zeta_{1}}{(\sin z\cos z)^{1/2}}.$
Note that by the scaling $z\to\epsilon z$, $\psi_{1}\to-\psi_{1}/(2\epsilon^{3/2})$ and change of parameter $\zeta_{1}=-2(1+\epsilon^{2}t_{1})$, in the $\epsilon\to 0$ limit, we obtain $\psi_{1}=(z^{2}+t_{1})/z^{1/2}$. This is $\psi_{1}$ (see (56)) of the even bi-spectral family of section 6.
In general, the terminating sequences of section 6 are the rational (times $z^{1/2}$) limits of the trigonometric sequences of the case $a=b=m+1/2$
$\psi_{m}={\cal W}[\varkappa_{m}],\quad\psi_{m+1}=\frac{{\cal W}[\varkappa_{m},\varkappa_{m+1}]}{{\cal W}[\varkappa_{m}]},\dots,\psi_{2m-1}=\frac{{\cal W}[\varkappa_{m},\varkappa_{m+1},\dots,\varkappa_{2m-2},\varkappa_{2m-1}]}{{\cal W}[\varkappa_{m},\varkappa_{m+1},\dots,\varkappa_{2m-2}]},$ (86)
where
$\varkappa_{m+i}=(\sin z\cos z)^{1/2-m}P^{(-m,-m)}_{m+i}(\cos 2z,\zeta_{m+i}),\quad i=0,1,\dots,m-1$ (87)
Here, polynomials $P^{(-m,-m)}_{n}$ are the special $l=m$ case of the para-Jacobi polynomials (“para-Gegenbauer” polynomials) [^15] They satisfy equation $\left((1-x^{2})\partial_{x}^{2}+2(m-1)x\partial_{x}+n(n-2m+1)\right)P^{(-m,-m)}_{n}(x,\zeta)=0$ which reduces to the standard hypergeometric form by the change of variable $w=(1-x)/2$.
$\begin{array}[]{l}P^{(-m,-m)}_{n}(x,\zeta)=\frac{(-2)^{n}(n-m)!n!}{(2n-2m)!}\sum_{k=0}^{n-m}\frac{(-1)^{n-k}(2n-2m-k)!}{k!(n-m-k)!(n-k)!}\left(\frac{1+x}{2}\right)^{n-k}+\\ \qquad\qquad\qquad\qquad\zeta\frac{(-2)^{n}(2n-2m+1)!(2m-n-1)!}{(n-m)!}\sum_{2(n-m)+1}^{n}\frac{(-1)^{n-k}(k-n+m-1)!}{k!(k+2m-2n-1)!(n-k)!}\left(\frac{1+x}{2}\right)^{n-k}\end{array}$
When we re-scale $z\to\epsilon z$, the eigenvalues scale as $\lambda_{i}\to\epsilon^{2}\lambda_{i}$, and in the rational limit we have confluent eigenvalues, all going to $\lambda_{i}\to\eta=0$. According to section 4 (see eqs. (42), (43)) in this limit the Wronskian representation of the Darboux transforms becomes
$\psi_{m+i}=\frac{\tau_{i+1}}{\tau_{i}},\quad\tau_{i}={\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{i}],\quad\tau_{0}=1,$ (88)
where now
$H_{0}{\cal X}_{1}(z)=0,\quad H_{0}{\cal X}_{i+1}(z)=c_{i}{\cal X}_{i}(z),\quad i=1\dots m-1.$
Here $H_{0}=-\partial_{z}^{2}+\frac{m^{2}-1/4}{z^{2}}$ is the rational limit of the trigonometric Schrodinger operator (83) with $a=b=m+1/2$, obtained by the scaling $z\to\epsilon z$, $H_{0}\to\epsilon^{-2}H_{0}$. In other words, ${\cal X}_{i}(z)$ are Laurent polynomials in $z^{1/2}$ that solve the chain
${\cal X}_{1}=z^{1/2-m}(z^{2m}+t_{m}),\quad-{\cal X}_{i+1}^{\prime\prime}+\frac{m^{2}-1/4}{z^{2}}{\cal X}_{i+1}=c_{i}{\cal X}_{i},\quad i=1,2,\dots,m-1,$ (89)
where $c_{i}$ are arbitrary constants. Although two constants of integration appear at each step of the chain, only one of them is essential. The chain terminates at the $(m-1)$th step, since the logarithmic term appears at the $m$th step. Function ${\cal X}_{i}$ is the $\epsilon\to 0$ limit of a linear combinations of $i$ first eigenfunctions (87):
$\epsilon^{3/2-2i-m}\varkappa_{m}(\epsilon z,\zeta_{m}),\quad\epsilon^{3/2-2i-m}\varkappa_{m+1}(\epsilon z,\zeta_{m+1}),\dots\epsilon^{3/2-2i-m}\varkappa_{m+i-1}(\epsilon z,\zeta_{m+i}),$
where the $(j-1)$th free parameter $\zeta_{m+j}$ is a linear combination of
$1,\quad\epsilon^{2m}t_{m},\quad\epsilon^{2m+2}t_{m+1},\quad\epsilon^{2m+4}t_{m+1},\dots\epsilon^{2(m+j-1)}t_{m+j}.$
Coefficients of linear combinations are constants determined by the condition of existence of the limit.
For example, in $m=2$ case, corresponding to the terminating sequence (57), we have
$\begin{array}[]{l}{\cal X}_{1}=\lim_{\epsilon\to 0}\epsilon^{-5/2}\varkappa_{2}(\epsilon z,2-2\epsilon^{4}t_{2}),\\ {\cal X}_{2}=\lim_{\epsilon\to 0}\epsilon^{-9/2}\left(\varkappa_{3}\left(\epsilon z,-\frac{1}{2}-\frac{3}{2}\epsilon^{4}t_{2}-\frac{1}{2}\epsilon^{6}t_{3}\right)-3\varkappa_{2}(\epsilon z,2-2\epsilon^{4}t_{2})\right).\end{array}$
Concluding this section we note that, in the trigonometric case, sequences of the static periodic configurations (86) can be continued with the help of solutions involving the ordinary Jacobi polynomials (see (85)). In other words, in difference from the rational limit, trigonometric sequences are not terminating. However, similarly to the rational case, these sequences depend only on finite numbers of free continuous parameters.
12 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in section 3, all possible static configurations of finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In the present section we will demonstrate that sequences of soliton $\tau$-functions (78,79) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To do this, we will introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with static and translating configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
In more detail, we consider the imaginary-time Schrodinger equation with time-dependent potential
$\frac{\partial\psi(z,t)}{\partial t}=H\psi(z,t),\quad H=-\partial_{z}^{2}+u(z,t).$ (90)
Here, $H$ can be presented in the form
$H=A^{*}A+(\log\varkappa)_{t}\,,$ (91)
where subscript $t$ stands for time derivative. In eq. (91), $\varkappa=\varkappa(z,t)$ is one of solutions of (90). The first-order differential operators $A$, $A^{*}$ are of the same form (29) as those involved into the time-independent Darboux transformation. Permuting factors in (91), we get the new Schrodinger operator
$\hat{H}=AA^{*}+(\log\varkappa)_{t}.$
Multiplying (91) by $A$ from the left, we see that $AH=\hat{H}A-A_{t}$, and therefore $\hat{\psi}=A\psi$ solves the new equation
$\frac{\partial\hat{\psi}}{\partial t}=\hat{H}\hat{\psi},\quad\hat{H}=-\partial_{z}^{2}+\hat{u}(z,t).$
The new potential $\hat{u}(x,t)$ and solution $\hat{\psi}$ are given by the same formulae as in the case of the time-independent Darboux transformations, i.e.
$\hat{\psi}=A\psi=\psi^{\prime}-(\log\varkappa)^{\prime}\psi,\quad\hat{u}=u-2(\log\varkappa)^{\prime\prime}.$ (92)
Now, we take the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. The corresponding imaginary-time Schrodinger (or inverse heat) equation admits polynomial solutions $p(z,t)$
$p_{t}=-p^{\prime\prime},\quad p=\prod_{i=1}^{l}(z-z_{i}(t)),$
where $z_{i}(t)$ obey the dynamical equations
$\frac{dz_{i}}{dt}=\sum_{j=1,j\not=i}^{l}\frac{2}{z_{i}-z_{j}},\quad i=1\dots l.$
Using $p$ as a seed function, $\varkappa=p$ in (92), we Darboux transform another polynomial solution $\psi$, $\psi_{t}=-\psi^{\prime\prime}$ into the rational solution
$\hat{\psi}=\frac{q}{p},\quad q=\psi^{\prime}p-p^{\prime}\psi.$ (93)
Then, the Schrodinger operator acquires the potential $\hat{u}=-2(\log p)^{\prime\prime}$, and the new equation $\hat{H}\hat{\psi}=\hat{\psi}_{t}$ writes, in terms of $p$ and $q$, as
$\left(-\partial_{z}^{2}-2(\log p)^{\prime\prime}\right)\left[q/p\right]=\left[q/p\right]_{t},$
or equivalently
$qp_{t}-pq_{t}=q^{\prime\prime}p-2p^{\prime}q^{\prime}+p^{\prime\prime}q.$ (94)
Thus, we incorporated dynamics to the Tkachenko equation (9). It is not difficult to see that further iterations of Darboux transformations will lead to equations of the same form for consecutive transforms of polynomials.
Equation (94) results in the following dynamical system for roots of $p$ and $q$, denoted as $z_{1},\dots,z_{l}$ and $z_{l+1},\dots,z_{l+m}$ respectively:
$\frac{dz_{i}}{dt}=2\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}},\quad i=1\dots N,\quad N=l+m,$ (95)
where negative and positive charges of equal magnitude ${\cal Q}_{i}=\pm 1$ are attached to roots of $q$ and $p$ correspondingly. If instead of polynomials $p$ and $\psi$ in (93) we take $pe^{kz-k^{2}t}$ and $\psi e^{kz-k^{2}t}$, we will get the Galilean transform of (95)
$\frac{dz_{i}}{dt}=2k+2\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}.$ (96)
Although this system seems to resemble that of moving vortices (4), in fact, it is quite different: In contrast to (4), the system (96) contains $dz_{i}/dt$ rather than complex conjugates $d\bar{z}_{i}/dt$. As a consequence, a systems of $N$ vortices is a hamiltonian system of dimension $2N$, while system of $N$ particles (96) has dimension $N$ and is not hamiltonian. Eq. (96) has several (quite artificial) physical interpretations. For example, we can interpret it as system of sinks/sources moving with the flow of the 2d ideal fluid. This and other interpretations make sense only when all $z_{i}$ move along the same line (e.g. all $z_{i}$ are real), except cases when all particles move with equal velocities, i.e. at fixed points of (96). These fixed points coincide with translating and static configurations (5,6) of systems of vortices.
Dynamics can be also incorporated into the bi-linear equation (8) for arbitrary $\Lambda$ as
$qp_{t}-\Lambda pq_{t}=q^{\prime\prime}p-2\Lambda p^{\prime}q^{\prime}+\Lambda^{2}p^{\prime\prime}q.$
The corresponding system has the form (95), where now ${\cal Q}_{i}=-1$ for roots of $p$, and ${\cal Q}_{i}=\Lambda$ for roots of $q$.
System (95), with arbitrary ${\cal Q}_{i}$, can be embedded into a hamiltonian system of newtonian particles interacting pairwise through inverse square potentials: Taking time derivative of eq. (95), then eliminating first order time derivatives by substituting them from (95), we get
$\frac{d^{2}z_{i}}{dt^{2}}=-4\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}({\cal Q}_{i}+{\cal Q}_{j})}{(z_{i}-z_{j})^{3}}$ (97)
(for details see e.g. [25] or [35] and references therein). For $\Lambda=1$, the sum of two charges vanishes ${\cal Q}_{i}+{\cal Q}_{j}=0$, when $i$ and $j$ belong to different species, and (97) decouples into two non-interacting sub-systems
$\frac{d^{2}z_{i}}{dt^{2}}=-\sum_{j:j\not=i}\frac{8}{(z_{i}-z_{j})^{3}},$
one consisting of charges ${\cal Q}_{i}=-1$ at $z_{1},\dots,z_{l}$ and another with ${\cal Q}_{i}=1$ at $z_{l+1},\dots,z_{l+m}$. These are completely integrable Calogero-Moser systems [30].
There is another case of decoupling, this time for three species of symmetrically situated particles: One particle of strength $a$ at the origin $z=0$, $2l$ particles of strength $-1$ at ±$\pm z_{1},\dots,\pm z_{l}$ and $2m$ particles of strength $1$ at ±$\pm z_{l+1},\dots,\pm z_{l+m}$. Here, due to the symmetry, we have reduction from $12(l+m)+$ 1 to $l+m$ degrees of freedom and decoupling into the two BC-type Calogero-Moser systems
$\frac{d^{2}z_{i}}{dt^{2}}=\frac{1-(2a\pm 1)^{2}}{z_{i}^{3}}-\sum_{j:j\not=i}\left(\frac{8}{(z_{i}-z_{j})^{3}}+\frac{8}{(z_{i}+z_{j})^{3}}\right)$
of $l$ and $m$ particles respectively (for review on Calogero-Moser systems related to different root systems see e.g. [34]). This case corresponds to the Darboux-transformed imaginary-time Schrodinger equation with
$H_{0}=-\partial_{z}^{2}+\frac{a(a-1)}{z^{2}}.$
Initial equation $H_{0}\psi=\psi_{t}$ has quasi-polynomial solutions $\psi=z^{a}\prod_{i=1}^{l}(z^{2}-z_{i}(t)^{2})$, and the Darboux transforms constructed from these solutions are quasi-rational. When $a$ is half-integer, fixed points of system (95) are configurations of even bi-spectral family from section 6.
Let us now return to the main topic of this section and consider equilibrium configurations of vortices. For these configurations $d^{2}z_{i}/dt^{2}=0$. Then form (97) it follows that the equilibrium conditions (6) imply the following $N$ locus conditions
$\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}({\cal Q}_{i}+{\cal Q}_{j})}{(z_{i}-z_{j})^{3}}=0,\quad i=1\dots N,$
i.e. the locus conditions are necessary conditions of equilibrium. In the case of two species with ${\cal Q}_{i}=\pm 1$, locus conditions decouple, and we have the separate set of locus conditions [^16] It is worth mentioning, that the theory of algebraically integrable systems also considers loci with multiple roots, where more general locus conditions are imposed [13, 12]. for each specie
$\sum_{j:j\not=i}\frac{1}{(z_{i}-z_{j})^{3}}=0,$ (98)
$l$ conditions for the first and $m$ conditions for the second specie respectively. It turned out [2] [^17] Conditions (98) first appeared in this famous paper by Airault, McKean and Moser, who introduced the term “locus”. that set of solutions of (98) is not empty only if the number of particles the locus is formed of is a triangular number $n(n+1)/2$. Then, it was shown [1] that for this number, roots of the Adler-Moser polynomials $P_{n}$ constitute complete solution of (98). As we saw in section 3, in systems of vortices in static equilibrium $l=n(n-1)/2$ and $m=n(n+1)/2$ particles form two loci that are roots of two consecutive Adler-Moser polynomials $P_{n-1}(z,t_{1},\dots,t_{n-2})$ and $P_{n}(z,t_{1},\dots,t_{n-1})$.
Conditions (98) also must hold for each specie of translating configurations (5) with ${\cal Q}_{i}=\pm 1$. We recall that for translating configurations $\deg p=\deg q$, i.e. $l=m$ (see section 8). This means that when $k\not=0$, $p$ and $q$ are Adler-Moser polynomials of the same degree, but with different sets of parameters: If $p=P_{n}(z,t_{1},\dots,t_{n-1})$, then $q=P_{n}(z+s_{0},s_{1},\dots,s_{n-1})$, where $s_{0},\dots,s_{n-1}$ are functions of $t_{1},\dots,t_{n-1}$ and $k$.
We also recall that, according to section 8, the rational Baker-Akhieser function $\psi=C(k)(q/p)e^{kz}$ determines translating configuration. It is an eigenfunction of a Schrodinger operator that is obtained by a sequence of Darboux transformations from free Schrodiger operator, i.e. $H\psi=k^{2}\psi$, where $H=-\partial_{z}^{2}-2(\log p)^{\prime\prime}$. From this we get the bilinear relation for $p$ and $q$:
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}+2k(q^{\prime}p-p^{\prime}q)=0.$ (99)
This relation coincides with the Tkachenko equation for translating configurations. Set of its solutions contains all translating patterns with ${\cal Q}_{i}=\pm 1$. The Baker-Akhieser function exists for any locus of $p$ and is unique for a given locus (see e.g. [13, 12] and references therein). Since the locus conditions are necessary conditions of an equilibrium, we come to the conclusion that all possible translating configurations consisting of a finite number of vortices with ${\cal Q}_{i}=\pm 1$ are determined by the Baker-Akhieser functions and are given by (69).
In the trigonometric (periodic) case, for each of two species of opposite charges the locus conditions write as
$\sum_{j:j\not=i}\frac{\cos(z_{i}-z_{j})}{\sin^{3}(z_{i}-z_{j})}=0.$
It is known (see e.g. [13, 12] and references therein) that all trigonometric locus configurations are given by roots of soliton $\tau$-functions of the KdV hierarchy. Therefore any pair $p$ and $q$ that does not have common roots and solves the periodic Tkachenko equation
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}+\left(l-m\right)^{2}pq=0,\quad p=\prod_{i=1}^{l}\sin(z-z_{i}),\quad q=\prod_{i=1}^{m}\sin(z-z_{l+i})$ (100)
consists of soliton $\tau$-functions. These functions are obtained with the help of factorization chain (34) starting from the free Schrodinger operator. In this chain of transformations
$H_{n}\psi^{(n)}=\eta_{n+1}\psi^{(n)},\quad H_{n}=-\partial_{z}^{2}-2(\log\tau_{n})^{\prime\prime},\quad\psi^{(n)}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{0}=1,$
i.e.
$\left(-\partial_{z}^{2}-2(\log\tau_{n})^{\prime\prime}\right)\left[\frac{\tau_{n+1}}{\tau_{n}}\right]=\eta_{n+1}\frac{\tau_{n+1}}{\tau_{n}}.$ (101)
Since degree of the leading term of the Laurent expansion of $\tau_{n+1}/\tau_{n}$ in $\exp({\rm i}z)$ equals $d_{n+1}-d_{n}$, where $d_{n}=\deg\tau_{n}$, from (101) we conclude that $\eta_{n+1}=(d_{n+1}-d_{n})^{2}$. Then, rewriting (101) in the bi-linear form, we get the Darboux chain for $\tau_{n}$:
$\tau_{n+1}^{\prime\prime}\tau_{n}-2\tau_{n+1}^{\prime}\tau_{n}^{\prime}+\tau_{n+1}\tau_{n}^{\prime\prime}+(d_{n+1}-d_{n})^{2}\tau_{n+1}\tau_{n}=0,\quad d_{n}=\deg\tau_{n},\quad\tau_{0}=1.$ (102)
The periodic analog of the Tkachenko equation (100) is of the same form as the Darboux chain equation (102). Then, taking the above fact and the fact that $p$ and $q$ are necessarily soliton $\tau$-functions (as follows from locus conditions) into account, we come to the conclusion that $p$ and $q$ are soliton $\tau$-functions related by Darboux transformation. Therefore, all possible static street patterns consisting of vortices of two species of opposite strengths are given by (78,79).
For translating streets we use arguments similar to those of the rational case and come to the conclusion that the complete set of translating street patterns with ${\cal Q}_{i}=\pm 1$ is determined by (78,82). It is worth mentioning that from the soliton theory it follows that here the relationship between $p$ and $q$ holds
$q=\tau_{n}(z,\sigma_{1},\dots,\sigma_{n}),\quad p=\tau_{n}(z,\zeta_{1},\dots,\zeta_{n}),\quad(k+{\rm i}k_{j})e^{2{\rm i}\sigma_{j}}=(k-{\rm i}k_{j})e^{2{\rm i}\zeta_{j}},$
where $\tau_{n}$ is given by (78,79). Without going into much details, we note that the above relationship can be derived using the Sato formula for the Baker-Akhieser function of the KP hierarchy and the correspondence between Hirota and Wronskian representations of the $n$-soliton $\tau$-functions (for introduction to theory of solitons see e.g. [39]).
Closing this section, we note that, in contrast to the cases considered above, the $\Lambda=2$ case is not exceptional as no decoupling into sub-systems of different species takes place there.
13 Conclusions and Open Problems
In the present paper we reviewed developments in application of factorization methods to the theory of point vortex patterns that took place during several last decades. In the $\Lambda=1$ case, these have connections with the theory of KdV hierarchy of integrable PDEs and the bi-spectral problem. In the $\Lambda=2$ case, connections with integrable hierarchies have been found by Demina and Kudryashov [15] and further developed by us in [28]. In difference from the KdV hierarchy, the Lax operators of these hierarchies are differential operators of the third order  [^18] We recall that in $\Lambda=1$ case the Lax operator is of the second order: Lax operator of the KdV hierarchy is a Schrodinger operator.. Indeed, in the $\Lambda=2$ case the bilinear equation (8) for $q_{n}$ and $p_{n}$ can be rewritten in the form of the Schrodinger equation at zero energy level
$\left(-\partial_{z}^{2}+u\right)\phi=0,\quad u=-6(\log q_{n})^{\prime\prime},$ (103)
where $\phi=p_{n}/q_{n}^{2}$ or $\phi=p_{n-1}/q_{n}^{2}$. According to (24), we can replace $p_{n}$ with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and get the following third-order differential equation for $\psi(z)=q_{n\pm 1}(z)/q_{n}(z)$
$L\psi=0,\quad L=\partial_{z}^{3}-u\partial_{z}.$
The third-order operator $L$ is a Lax operator of the Sawada-Kotera hierarchy and $q_{n}$ is its polynomial $\tau$-function. Such polynomials can be constructed with help of the Darboux transformations at “zero level” [28]
[…]
where primitives include arbitrary constants of integration. Corresponding third-order operators can be obtained by applications of sequence of the above transformations to $\partial_{z}^{3}$, while $q_{n}$ can be expressed in form of a Pfaffian [20].
Here we just note that, since $L$ is not a Schrodinger operator, the Baker-Akhiezer function $\psi(k,z)=P_{n}(k,z)/q_{n}(z)e^{kz}$ of the Sawada-Kotera hierarchy
$L\psi(k,z)=k^{3}\psi(k,z)$
does not correspond to stationary configurations when $k\not=0$. Also, soliton solutions of this hierarchy do not correspond to stationary configurations on the cylinder. Only when $k=0$, the above eigenvalue problem reduces to solution of the Schrodinger equation (103) with $\phi(z)=\lim_{k\to 0}k^{N_{n}}\frac{\partial\psi(k,z)}{\partial z}$ corresponding to static configurations. Here factor $k^{N_{n}}$, with suitably chosen $N_{n}$, insures finiteness of the limit.
In difference from the $\Lambda=1$ case, the $\Lambda=2$ case seems not to have as many exceptional properties: In particular, the fact that the soliton Baker-Akhiezer functions of the Sawada-Kotera hierarchy are not eigenfunctions of the Schrodinger operators suggests absence of non-trivial non-terminating sequences on the cylinder in the $\Lambda=2$ case.
We searched for sequences of configuration on the cylinder by direct solution of the trigonometric analog of the bilinear equation (8)
$\tau_{n-1}^{\prime\prime}\tau_{n}-2\gamma_{n}\tau_{n-1}^{\prime}\tau_{n}^{\prime}+\gamma_{n}^{2}\tau_{n-1}\tau_{n}^{\prime\prime}+(d_{n-1}-\gamma_{n}d_{n})^{2}\tau_{n-1}\tau_{n}=0,\quad\gamma_{n}\gamma_{n+1}=1,$ (104)
and in the $\Lambda=2$ case we found only short terminating sequences, for example
$\tau_{0}=\xi-\frac{1}{\xi},\tau_{1}=\xi^{6}-4\xi^{4}+5\xi^{2}+\frac{5t_{1}}{\xi^{2}}-\frac{4t_{1}}{\xi^{4}}+\frac{t_{1}}{\xi^{6}},\,\tau_{2}=\xi^{5}-5\xi^{3}-t_{2}\xi+\frac{t_{2}}{\xi}-\frac{5t_{1}}{\xi^{3}}+\frac{t_{1}}{\xi^{5}},$
where $\xi=e^{{\rm i}z}$ and $\gamma_{0}=2$. The above example can be continued one more step if we impose condition $t_{1}=t_{2}^{2}/5$. Then we will get another terminating sequence with $\tau_{1}=\tau_{1}(\xi,t_{2})$, $\tau_{2}=\tau_{2}(\xi,t_{2})$ and $\tau_{3}=\tau_{3}(\xi,t_{2},t_{3})$ , where $\deg\tau_{3}=14$.
Without going into much detail we just note that in such type of sequences $\tau_{1}$ is a solution of an ODE that is reducible to the hypergeometric form. Then we can apply KWCC transformations (60) to $\psi_{i}=\tau_{i}^{\gamma_{i}}/\tau_{i-1}$, $i=1,2$, to continue the sequence. The transformation can be applied at most twice. Terminating configurations obtained in this way can depend on no more than two non-trivial free parameters (they depend on two parameters when $\deg\tau_{1}=6,8,10,\dots$). Here, a natural question arises: If the above procedure generates all possible static vortex streets patterns of two distinct species of vortices with $Q_{i}\in\{-1,2\}$
Regarding the static patterns formed by finite number of vortices, we may ask the question: Whether the main sequences (23), (25), (26) for the $\Lambda=1$ and $\Lambda=2$ cases are the only possible non-trivial non-terminating sequences of the KWCC transforms. Are there non-trivial sequences (60) with $\gamma_{i}$ different from (61,62) that do not terminate?
Finally, classification of all stationary vortex patterns and understanding the role of factorization methods in this classification are ultimate questions to address in this subject.
Concluding this review, we would like to mention related problem from the theory of algebraically integrable systems, because our interest in vortex patterns stems from that theory: In the case of odd family of periodic configurations, it is convenient to rewrite the bi-linear chain (102) in terms of homogenous polynomials in two variables, replacing $\tau_{n}(z)=\prod_{i=1}^{d_{n}}\sin(z-z_{i})$ with
$\tau_{n}(X,Y)=R^{d_{n}}\tau_{n}(z)=\prod_{i=1}^{d_{n}}(Y\cos z_{i}-X\sin z_{i}),$
where
$X=R\cos z,\quad Y=R\sin z.$
In the new variables, the chain (102) takes the following form
$\tau_{n}\Delta\tau_{n+1}-2(\nabla\tau_{n+1}\cdot\nabla\tau_{n})+\tau_{n+1}\Delta\tau_{n}=0,\quad\tau_{0}=1,$ (105)
where $\Delta$ and $\nabla$ stand for two-dimensional laplacian and gradient respectively. This chain is called harmonic [5], because at the first step $\Delta\tau_{1}=0$, i.e. $\tau_{1}$ is a harmonic function, and $\psi_{n}=\tau_{n+1}/\tau_{n}$ are natural generalizations of harmonic functions [^19] For example, $\psi_{n}$ satisfy quadrature identities in “harmonic” quadrature domains (see e.g. [26, 27] and references therein).. The harmonic chain can be generalized to any number of dimensions: Non-terminating solutions of (105) are $\tau$-functions for potentials of algebraically integrable Schrodinger operators (see e.g. [5], [12]). One can view (105) as a generalization of the Darboux chain (34), that starts from the free Schrodinger operator [^20] Chain that starts from $u_{0}\not=0$ has the form $\tau_{n}\Delta\tau_{n+1}-2(\nabla\tau_{n+1}\cdot\nabla\tau_{n})+\tau_{n+1}\Delta\tau_{n}-u_{0}\tau_{n}\tau_{n+1}=0$, $\tau_{0}=1$., to any number of dimensions. In one dimension, the Adler-Moser polynomials constitute complete set of non-terminating solutions of the chain. In two dimensions, the soliton related solutions exhaust all non-terminating solutions in the class of homogeneous polynomials: Family of two-dimensional solutions consists of an infinite number of branches labelled by $0<k_{1}<k_{2}<\dots$, rather than a single sequence in one dimension. Complete classification of non-terminating solutions in all dimensions is an open hard problem (see e.g. [5, 12]) [^21] It is conjectured that families listed in [12] constitute complete solution of the problem..